OpenStudy (anonymous):

Find the derivative of f(x) = 12x^2 + 8x at x = 9 256 -243 288 224

OpenStudy (sidsiddhartha):

use the chain rule \[\Large \frac{ d }{ dx }(x^n)=n*x^{n-1}\]

OpenStudy (sidsiddhartha):

can u do it now?

OpenStudy (anonymous):

Never learned a chain rule

OpenStudy (sidsiddhartha):

ok just use the above formula \[u ~~have ~~~f(x)=12x^2+8x\] now using the formula \[f'(x)=12[2.x^{2-1}]+8=12.[2x]=8=24x+8\] now just put x=9 \[f'(x=9)=24*9+8=224\]

OpenStudy (anonymous):

Alright, I got it now. Thanks, man!

OpenStudy (sidsiddhartha):

yw bud !!